a jar of instant coffee contains 16 and one third ounces of coffee. how many cups of coffee containing 1 and one sixth ounces of coffee can be made?
Answer by lwsshak3(6770)

a jar of instant coffee contains 16 and one third ounces of coffee. how many cups of coffee containing 1 and one sixth ounces of coffee can be made?
**
(16+1/3)/(1+1/6)
=(49/3)/(7/6)
=49/3*6/7
=7*2
=14
cups of coffee containing 1 and one sixth ounces of coffee that can be made=14